Calculate the molality of a solution where 0.5 grams of toluene (C7H8) is dissolved in 225 grams of (C6H6). Molecular weight = C7H8=7×12+1×8=92 grams/mole
1
Expert's answer
2020-05-03T15:23:56-0400

Solution.

The molality is calculated by: "M = \dfrac{\nu_{solute}}{m_{solvent}};"

"\nu(C_7H_8) = \dfrac{m(C_7H_8)}{M(C_7H_8)};"

"M(C_7H_8) = 92g\/mol;"

"\nu(C_7H_8) = \dfrac{0.5g}{92g/mol}=0.0054mol;"

 "\nu_{solute}" is the number of mole solute, and "m_{solvent}"  is the mass of the solvent in kg.

"m_{solvent}=225g=0.225kg;"

The molality then: "M = \dfrac{0.0054mol}{0.255kg}=0.024mol/kg;"

Answer: M = 0.024mol/kg.